ten years ago mike was four times old as lisa, now he is twice as old as liza, how old are they now?
Answer by ikleyn(52787)

.

If x is Lisa's present age, then Mike's present age is 2x years.


Ten years ago . . . 


    4*(x - 10) = 2x-10


Simlify and solve


    4x - 40 = 2x - 10

    4x - 2x = 40- 10

    2x      = 30

     x      = 30/2 = 15.


ANSWER.  Lisa is 15 years old.  Mike is 2*15 = 30 years old.